Solve the following differential equation

dx/dt+dy/dt-2x-4y=et

dx/dt+dy/dt-y=e4t




1
Expert's answer
2021-08-10T13:13:09-0400
"2x+4y+e^{t}=y+e^{4t}"

"x=-\dfrac{3}{2}y-\dfrac{1}{2}e^{t}+\dfrac{1}{2}e^{4t}"

"x'=-\dfrac{3}{2}y'-\dfrac{1}{2}e^{t}+2e^{4t}"

"-\dfrac{3}{2}y'-\dfrac{1}{2}e^{t}+2e^{4t}+y'+3y+e^{t}-e^{4t}-4y=e^{t}"

"\dfrac{1}{2}y'+y=-\dfrac{1}{2}e^{t}+e^{4t}"

"y'+2y=-e^{t}+2e^{4t}"


"\mu(t)=e^{2t}"

"e^{2t}y'+2e^{2t}y=-e^{3t}+2e^{6t}"

"(e^{2t}y)'=-e^{3t}+2e^{6t}"

Integrate


"\int d(e^{2t}y)=\int(-e^{3t}+2e^{6t})dt"

"e^{2t}y=-\dfrac{1}{3}e^{3t}+\dfrac{1}{3}e^{6t}+C_1"

"y(t)=-\dfrac{1}{3}e^{t}+\dfrac{1}{3}e^{4t}+C_1e^{-2t}"

"x(t)=\dfrac{1}{2}e^{t}-\dfrac{1}{2}e^{4t}-\dfrac{3C_1}{2}e^{-2t}-\dfrac{1}{2}e^{t}+\dfrac{1}{2}e^{4t}"




"x(t)=-\dfrac{3C_1}{2}e^{-2t}"

"y(t)=-\dfrac{1}{3}e^{t}+\dfrac{1}{3}e^{4t}+C_1e^{-2t}"